# Design and Implementation of Energy-Efficient Wireless Tire Sensing System with Delay Analysis for Intelligent Vehicles

### v2

@@ Abstract
The growing prevalence of Internet of Things (IoT) technologies has led to a rise in the popularity of intelligent vehicles that incorporate a range of sensors to monitor various aspects, such as driving speed, fuel usage, distance proximity and tire anomalies. Nowadays, real-time tire sensing systems play important roles for intelligent vehicles in increasing mileage, reducing fuel consumption, improving driving safety, and reducing the potential for traffic accidents. However, the current tire sensing system drains a significant vehicle’ energy and lacks effective collection of sensing data, which may not guarantee the immediacy of driving safety. Thus, this paper designs an energy-efficient wireless tire sensing system (WTSS), which leverages energy-saving techniques to significantly reduce power consumption while ensuring data retrieval delays during real-time monitoring. Additionally, we mathematically analyze the worst-case transmission delay and sensor reception ratio of the system to ensure the immediacy based on the collision probabilities of sensor transmissions. This system has been implemented and verified by the simulation and field train experiments. These results show that the proposed scheme provides enhanced performance in energy efficiency up to 76.5% in average and identifies the worst transmission delay accurately.
@@ I Introduction
To address the above issue, we propose an energy-efficient IoT-enabled wireless tire sensing system (WTSS) that incorporates wireless sensors and multiple RF transmission modules to enable real-time monitoring. We develop and demonstrate this system based on wireless pressure and temperature sensors, which are widely used in intelligent vehicles [16]. These sensors are low-end and typically comprise a radio-frequency (RF) transmitter module, sensor cells, and a unique sensor ID, which make it possible for the receiver end to decode the pressure and temperature in real time [17]. In addition, an energy-saving mechanism is designed at the system receiver to conserve energy. This mechanism enforces the receiver to perform wake-up and sleep operations to maintain system reliability and immediacy. Specifically, an analytic scheme is developed to identify the wake-up and sleep schedule and also analyze the sensing delays to achieve real-time identification and avoid potential consequences. The extended simulations and experiments of the field trial will show that the proposed system can save up to $76.5\%$ of energy on average while ensuring the worst transmission delay.
@@ IV-A Analytic Scheme
$\frac{(n+1)+(C^{\text{min}}_{L}\cdot C_{L})}{[1+S]}\in\mathbb{N},$
@@
According to Corollary 1.2.1, we know that for all $n=1,2,\dots,C_{L}$, their corresponding $C^{\text{min}}_{L}$ will not exceed $S$, i.e., $0\leq C^{\text{min}}_{L}\leq S$. In the following, we prove this lemma by contradiction. Assume it exists a slot $n^{\prime}\neq C_{L}-S$, which insures a delay worse than $n=C_{L}-S$. We consider this in two cases: one is $n^{\prime}<C_{L}-S$ and $n^{\prime}>C_{L}-S$ as follows.
@@
The probability of this total delay333 Note that if the system does not receive a sensor over $T_{delay}$ slots, it will re-activate such sensor to randomly transmit again to avoid such sensor being collided by the same pattern., denoted as $Pr(W_{delay}^{t})$, is determined by the probability of encountering a collision $Pr(c)$, raised to the power of $t$, multiplied by the probability of successful reception $Pr(s)$, and it is with a probability given by
@@ V-B Expected Worst Transmission Delay
Fig. 7 shows the expected worst transmission delay (in seconds) received in the system for different numbers of sensors transmitting, where $E[W_{delay}^{total}]=\sum_{t=0}^{\infty}Pr(c)\times t\times W_{delay}+Pr(s)\times W_{delay}$. Based on different numbers of transmitting sensors in the system, it is shown that the expected delay of Real-time Experiment is always higher than Mathematical Analysis. This is due to a higher collision probability $(Pr(c))$ with an increasing number of transmitting sensors, resulting in a higher expected worst transmission delay.
@@ V-C Average Transmission Delay
V-D Successful Sensor Receiving Rate
[…]
Figure 9: Comparisons on successful sensor reception rate
V-E Power Saving Ratio
@@
Applications Sensors Wireless Channel Trans. Interval    (in ms) Delay Constraints (in ms) Energy Consumption *Suggested Parameters
@@
V-F Cumulative Correctness Rate (CCR)
Finally, the correctness rate, the ratio of the theoretical values derived from the Mathematical Analysis (Fig. 11) to the corresponding Real-time Experiment values for Worst Transmission Delay, Average Transmission Delay, and Successful Sensor Receiving Rate. CCR is defined as:
@@ V-F Cumulative Correctness Rate (CCR)
V-G Discussion of Comparison and Analysis with Other Applications

### v3

@@ Abstract
The growing prevalence of Internet of Things (IoT) technologies has led to a rise in the popularity of intelligent vehicles that incorporate a range of sensors to monitor various aspects, such as driving speed, fuel usage, distance proximity and tire anomalies. Nowadays, real-time tire sensing systems play important roles for intelligent vehicles in increasing mileage, reducing fuel consumption, improving driving safety, and reducing the potential for traffic accidents. However, the current tire sensing system drains a significant vehicle’ energy and lacks effective collection of sensing data, which may not guarantee the immediacy of driving safety. Thus, this paper designs an energy-efficient wireless tire sensing system (WTSS), which leverages energy-saving techniques to significantly reduce power consumption while ensuring data retrieval delays during real-time monitoring. Additionally, we mathematically analyze the worst-case transmission delay of the system to ensure the immediacy based on the collision probabilities of sensor transmissions. This system has been implemented and verified by the simulation and field trial experiments. These results show that the proposed scheme provides enhanced performance in energy efficiency and accurately identifies the worst transmission delay.
@@ I Introduction
To address the above issue, we propose an energy-efficient IoT-enabled wireless tire sensing system (WTSS) that incorporates wireless sensors and multiple RF transmission modules to enable real-time monitoring. We develop and demonstrate this system based on wireless pressure and temperature sensors, which are widely used in intelligent vehicles [16]. These sensors are low-end and typically comprise a radio-frequency (RF) transmitter module, sensor cells, and a unique sensor ID, which make it possible for the receiver end to decode the pressure and temperature in real time [17]. In addition, an energy-saving mechanism is designed at the system receiver to conserve energy. This mechanism enforces the receiver to perform wake-up and sleep operations to maintain system reliability and immediacy. Specifically, an analytic scheme is developed to identify the wake-up and sleep schedule and also analyze the sensing delays to achieve real-time identification and avoid potential consequences. The extended simulations and experiments of the field trial will show that the proposed system can save energy while ensuring the worst transmission delay.
@@ IV-A Analytic Scheme
$\frac{(n-1)+(C^{\text{min}}_{L}\cdot C_{L})}{[1+S]}\in\mathbb{N},$
@@
According to Corollary 1.2.1, we know that for all $n=1,2,…,C_{L}$, their corresponding $C^{\text{min}}_{L}$ will not exceed $S$, i.e., $0\leq C^{\text{min}}_{L}\leq S$. In the following, we prove this lemma by contradiction. Assume it exists a slot $n^{\prime}\neq C_{L}-S$, which insures a delay worse than $n=C_{L}-S$. We consider this in two cases: one is $n^{\prime}<C_{L}-S$ and the other is $n^{\prime}>C_{L}-S$ as follows.
@@
The probability of this total delay333 Note that if the system does not receive a sensor over $T_{delay}$ slots, the system will re-activate such sensor to randomly transmit again to avoid such sensor being collided by the same pattern., denoted as $Pr(W_{delay}^{t})$, is determined by the probability of encountering a collision $Pr(c)$, raised to the power of $t$, multiplied by the probability of successful reception $Pr(s)$, and it is with a probability given by
@@ V-B Expected Worst Transmission Delay
Fig. 7 shows the expected worst transmission delay (in seconds) received in the system for different numbers of sensors transmitting, where $E[W_{delay}^{total}]=\sum_{t=0}^{\infty}(t+1)\times W_{delay}\cdot Pr(c)^{t}\times Pr(s)$. Based on different numbers of transmitting sensors in the system, it is shown that the expected delay of Real-time Experiment is always higher than Mathematical Analysis. This is due to a higher collision probability $(Pr(c))$ with an increasing number of transmitting sensors, resulting in a higher expected worst transmission delay.
@@ V-C Average Transmission Delay
V-D Power Saving Ratio
@@
Applications Sensors Wireless Channel Trans. Interval    (in ms) Delay Requirements (in ms) Energy Consumption *Suggested Parameters
@@
V-E Cumulative Correctness Rate (CCR)
Finally, the correctness rate, the ratio of the theoretical values derived from the Mathematical Analysis (Fig. 10) to the corresponding Real-time Experiment values for Worst Transmission Delay and Average Transmission Delay. CCR is defined as:
@@ V-E Cumulative Correctness Rate (CCR)
V-F Discussion of Comparison and Analysis with Other Applications